OpenStudy (anonymous):

Given: m<1=24 degrees, m<3=24 degrees, <1 and <2 are complementary, <3 and <4 are complementary Prove: <2 congruent <4

OpenStudy (mathstudent55):

Do you need a two-column proof?

OpenStudy (anonymous):

yeah its a seven step proof

OpenStudy (mathstudent55):

How do you know the number of steps?

OpenStudy (anonymous):

We were given the number of steps

OpenStudy (mathstudent55):

Ok. This problem deals with complementary angles. Do you know what complementary angles are?

OpenStudy (anonymous):

yes angles that add up to 90 degrees

OpenStudy (mathstudent55):

Great. The first step in writing a proof is to figure out the entire problem. Then you can write it all out.

OpenStudy (mathstudent55):

We are told this. Angles 1 and 3 both measure 24 degrees. They are congruent. Angle 1 is complementary to angle 2. Angle 3 is complementary to angle 4.

OpenStudy (mathstudent55):

Angles 2 and 4 are complementary to two congruent angles. That means angles 2 and 4 must congruent to each other.

OpenStudy (mathstudent55):

That is the reasoning behind this proof. Now we can actually write the proof.

OpenStudy (mathstudent55):

Do you usually write all the given in one statement?

OpenStudy (anonymous):

well in this problem i do not believe it is needed to split the given

OpenStudy (mathstudent55):

Right. We can have all the given as one statement at the beginning. Statements Reasons 1. m<1=24 degrees 1. Given m<3=24 degrees <1 and <2 are complementary <3 and <4 are complementary 2. m<1 + m<2 = 90 2. Definition of complementary angles 3. m<3 + m<4 = 90 2. Definition of complementary angles

OpenStudy (mathstudent55):

4. 24 + m<2 = 90 4. Substitution property of equality 5. 24 + m<4 = 90 5. Substitution property of equality 6. m<2 = 66 6. Subtraction property of equality 7. m<4 = 66 7. Subtraction property of equality 8. m<2 = m<4 8. Transitive property of equality

OpenStudy (mathstudent55):

9. <2 is congruent to <4 9. Definition of congruent angles.